OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. −4x + 6y = 20 4x − 3y = −2 −2x + 4y = 18 4x − 8y = −36 3x − 2y = 5 −6x − 4y = −10 5x − 3y = 12 15x − 9y = 24

OpenStudy (anonymous):

4x - 4y = -10

OpenStudy (trojanpoem):

D: 15x - 9y = 24 (divide over 3) 5x - 3y = 12 Same equations, that means we have insufficient number of equation to get a solely answer so we have infinite. You could use matrices, if you want I can provide a solution using them.

OpenStudy (trojanpoem):

Matrix ranks. Creamer rule. ( not sure of his name :) )

OpenStudy (anonymous):

THANKS @TrojanPoem

OpenStudy (trojanpoem):

You're welcome.